OpenStudy (anonymous):

Solve 18.2 = y - 11 for y.

OpenStudy (anonymous):

add 11 to both sides

OpenStudy (anonymous):

Solve -5w = -80 for w= what

OpenStudy (anonymous):

29.2

OpenStudy (anonymous):

w = 16

OpenStudy (anonymous):

mine or hers

OpenStudy (nirmalnema):

18.2 = y - 11 for y. 36=y-11 y=36+11 =47

OpenStudy (nirmalnema):

18.2 means 18*2=36